# Study on data analysis for Ives-Stilwell-type experiments

### v5

@@ Abstract
Ives-Stilwell experiment in 1938 was a historic experiment for confirming Einstein’s special relativity, and various modern types have been repeated by use of laser technology. In this paper, we reveal and solve a long-lasting unrecognized fundamental issue that the data analysis for all those experiments was not consistent with Einstein’s definition of the relativistic Doppler effect so that the Doppler effect was not actually confirmed. Based on existing data, we propose a justified data analysis and correctly confirm the Doppler effect in the Ives-Stilwell-type experiment, thus resulting in a great advance in the experimental verification of Lorentz invariance via the Doppler effect.
@@ I Introduction
The relativistic Doppler effect is one of the basic results of Einstein’s special relativity, and it states that a plane light wave or a single laser beam or a single photon in free space, observed in any inertial frames of reference, propagates at the same speed but it may exhibit different frequencies and aberrations r1 . The 1938 Ives-Stilwell experiment r2 was a historic research work for confirming Einstein’s special relativity via the Doppler effect, and various modern types of the experiment r3 ; r4 ; r5 ; r6 ; r7 have been repeated by use of laser technology. The concluding result from an Ives-Stilwell-type experiment is presented in the Letter r7 by an international group of collaborators after 15 years of efforts r8 . However the data analysis method for all those experiments r2 ; r3 ; r4 ; r5 ; r6 ; r7 was not consistent with Einstein’s definition of the relativistic Doppler effect so that the Doppler effect was not actually confirmed. Here, based on existing experimental data r7 , we propose a justified analysis method and correctly confirm the Doppler effect in the Ives-Stilwell-type experiment, thus successfully revealing and solving a long-lasting unrecognized fundamental issue in the experimental confirmation of special relativity.
II Definitions of measurement accuracies and analysis of experimental data
In the excellent Letter, Botermann et al. r7 report that the measurement accuracy of relativistic Doppler effect has reached the order of $10^{-9}$ in an Ives-Stilwell-type experiment, where a moving high-energy ion beam is set at resonance with two laser beams aligned parallel and antiparallel relative to the ion beam, and the ions with two very close transition frequencies behave as Doppler-shifted laser frequency detector.
@@ II Definitions of measurement accuracies and analysis of experimental data
where $\gamma=(1-\beta^{2})^{-1/2}$ is the relativistic factor. Note that either $\nu_{a\textrm{-sh}}$ or $\nu_{p\textrm{-sh}}$ can be taken as the true value of measured $\nu_{1}$ or $\nu_{2}$.
@@ III Conclusions and remarks
In this paper, we have revealed and solved a long-lasting unrecognized fundamental issue on data analysis for Ives-Stilwell-type experiments. We have shown that $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$, given by Eq. (1), is an expression of the coupling of various accuracies for two individual laser beams, and it cannot represent the accuracy for any single beam of them, namely it cannot represent the measurement accuracy of the Doppler effect defined by Einstein. To confirm this, by use of the experimental data provided in the Letter where the measured ion beam velocity is given by $\beta=0.338$ r7 , we calculated all four accuracies for single laser beams, showing that the accuracy for a single laser beam is much lower, in the order of $10^{-4}$, different from the accuracy reported in the Letter by five orders of magnitude.
@@
One might argue that although the measurement accuracy of the Doppler effect can only be represented by that from a single laser beam, it is possible to correlate the two measurements from antiparallel and parallel laser beams in order to improve accuracy without challenging the physical phenomenon. However this is not true, because the Doppler phenomenon is judged by checking the accuracy of Doppler effect from a single laser beam, and the correlation of the two measurements may change the accuracy, and further challenge the judgment of Doppler phenomenon.
@@
The present paper might be another interesting example, with the same experimental data used to form an alternative chain of evidence, leading to a unique and definite conclusion.
@@
Data availability. Data underlying the results presented in this paper are available within the paper and Supplement 1.
Supplemental document. See Supplement 1 for supporting content r13 .

### v11

@@ Abstract
Ives-Stilwell experiment in 1938 was a historic experiment for confirming Einstein’s special relativity, and various modern types have been repeated by use of laser technology. In this paper, we reveal and solve a long-lasting unrecognized fundamental issue that the data analysis for all those experiments was not consistent with Einstein’s definition of the relativistic Doppler effect so that the Doppler effect was not actually confirmed. For example, in the Letter [Phys. Rev. Lett. 113, 120405 (2014)] the definition of the measurement accuracy of Doppler effect, given by $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$, is not physical because Einstein’s Doppler formula cannot be confirmed even when $\varepsilon=0$ holds. Based on existing data, we propose a justified data analysis and correctly confirm the Doppler effect in the Ives-Stilwell-type experiment, thus resulting in a great advance in the experimental verification of Lorentz invariance via the Doppler effect. The proposed data analysis in the paper is quite straightforward, but it is the result of the correct way of thinking based on first principles: (i) Einstein’s Doppler effect refers to the same photon (laser beam) exhibiting different frequencies observed in two inertial frames of relative motion, and (ii) the quantity (measurement accuracy) used as a measure to confirm the effect must be able to confirm Einstein’s Doppler formula itself. However, the way of thinking set up in the 1938 Ives-Stilwell experiment was incorrect because it was not consistent with the first principles, although it had been followed by quite a few generations of physicists. Even Einstein seemed unaware of this subtle problem with such a far-reaching impact, which highlights the significance of the way of thinking provided in the present paper for analyzing and confirming Einstein’s Doppler effect.
@@ I Introduction
The relativistic Doppler effect is one of the basic results of Einstein’s special relativity, and it states that a plane light wave or a single laser beam or a single photon in free space, observed in any inertial frames of reference, propagates at the same speed but it may exhibit different frequencies and aberrations r1 . The 1938 Ives-Stilwell experiment r2 was a historic research work for confirming Einstein’s special relativity via the Doppler effect, and various modern types of the experiment r3 ; r4 ; r5 ; r6 ; r7 have been repeated by use of laser technology. The concluding result from an Ives-Stilwell-type experiment is presented in the Letter r7 by an international group of collaborators after 15 years of efforts r8 . However the data analysis method for all those experiments r2 ; r3 ; r4 ; r5 ; r6 ; r7 was not consistent with Einstein’s definition of the relativistic Doppler effect so that the Doppler effect was not actually confirmed. For example, the definition of the measurement accuracy of Doppler effect in r7 is not physical (see Appendix A). Here, based on existing experimental data r7 , we propose a justified analysis method and correctly confirm the Doppler effect in the Ives-Stilwell-type experiment, thus successfully revealing and solving a long-lasting unrecognized fundamental issue in the experimental confirmation of special relativity.
II Accuracies and data analysis
In the excellent Letter, Botermann et al. r7 report that the measurement accuracy of relativistic Doppler effect has reached the order of $10^{-9}$ in an Ives-Stilwell-type experiment, where a moving high-energy ion beam is set at resonance with two laser beams aligned parallel and antiparallel relative to the ion beam, and the ions with two very close transition frequencies behave as Doppler-shifted laser frequency detector, as shown in Fig. 1.
Figure 1: Schematic diagram of Ives-Stilwell-type experiment based on optical-optical double resonance (OODR) spectroscopy r7 . Two laser beams are aligned antiparallel $(\nu_{a})$ and parallel $(\nu_{p})$ relative to an ion beam moving at $\beta$. The ion beam has two transition frequencies $\nu_{1}$ and $\nu_{2}$, functioning as Doppler-shifted laser frequency detector.
@@ II Accuracies and data analysis
where $\gamma=(1-\beta^{2})^{-1/2}$ is the relativistic factor. Note that either $\nu_{a\textrm{-sh}}$ or $\nu_{p\textrm{-sh}}$ can be taken as the true value of measured $\nu_{1}$ or $\nu_{2}$. From Eqs. (3) and (4) it can be seen that the antiparallel and parallel laser beams are of different types of Doppler effects.
@@ III Conclusions and remarks
In this paper, we have revealed and solved a long-lasting unrecognized fundamental issue on data analysis for Ives-Stilwell-type experiments. We have shown that $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$, given by Eq. (1), is an expression of the coupling of various accuracies for two individual laser beams, and it cannot represent the accuracy for any single beam of them, namely it cannot represent the measurement accuracy of the Doppler effect defined by Einstein. To confirm this, by use of the experimental data provided in the Letter where the measured ion beam velocity is given by $\beta=0.338$ r7 , we calculated all four accuracies for single laser beams, showing that the accuracy for a single laser beam is much lower, in the order of $10^{-4}$, different from the accuracy reported in the Letter by five orders of magnitude; especially, we have proved that the quantity $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$ as the definition of the measurement accuracy of Doppler effect in r7 is unphysical, because Einstein’s Doppler formula cannot be confirmed even when $\varepsilon=0$ holds, as shown in Appendix A.
@@
One might argue that although the measurement accuracy of the Doppler effect can only be represented by that from a single laser beam, it is possible to correlate the two measurements from antiparallel and parallel laser beams, which are of different types of Doppler effects, in order to improve accuracy without challenging the physical phenomenon. However this is not true, because the Doppler phenomenon is judged by checking the accuracy of Doppler effect from a single laser beam, and the correlation of the two measurements may confuse the physics of Doppler effect and change the accuracy, and further challenge the judgment of Doppler phenomenon.
It should be reiterated that the essential flaw of the Letter r7 is the definition of the measurement accuracy for Doppler effect inconsistent with Einstein Doppler formula, instead of whether $\beta=0.338$ or $\beta=0.338,377,22$ should be taken in data analysis, although the correct $\beta=0.338$ is taken in the paper.
@@
The present paper might be another interesting example, with the same experimental data used to form an alternative chain of evidence, leading to a unique and definite conclusion. Although the chain of evidence for single laser beams is quite straightforward, it is the result of the correct way of thinking based on first principles: (i) Einstein’s Doppler effect refers to the same photon (laser beam) exhibiting different frequencies observed in two inertial frames of relative motion, and (ii) the quantity (measurement accuracy) used as a measure to confirm the effect must be able to confirm Einstein’s Doppler formula itself. However, the way of thinking set up in the 1938 Ives-Stilwell experiment r2 was incorrect because it was not consistent with the first principles, although it had been followed by quite a few generations of physicists r3 ; r4 ; r5 ; r6 ; r7 . Even Einstein seemed unaware of this subtle problem with such a far-reaching impact, which highlights the significance of the way of thinking provided in the present paper for analyzing and confirming Einstein’s Doppler effect.
@@
Acknowledgment. The author acknowledges the Letter by Botermann et al. r7 without which the present work would not have been possible. Note: Like citing published theoretical formulas (Einstein’s Doppler formula, for example), it is allowable and legitimate to cite the experimental data in r7 to support the theory developed in the present paper, and it is not in violation of any ethics guidelines.
@@
Data availability. Data underlying the results presented in this paper are available within the paper and Supplement 1 r13 .
Appendix A Einstein’s Doppler formula not confirmed at $\varepsilon=0$
In this appendix, by specific calculations based on existing experimental data, we show that the quantity $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$ as the definition of the measurement accuracy of Doppler effect in r7 is physically incorrect, because Einstein’s Doppler formula, $\nu_{2}=\gamma(1+\beta)\nu_{a}$, cannot be confirmed even when $\varepsilon=0$ holds.
For $\varepsilon=0$, from Eq. (1) and Eq. (11) we have $\nu_{a}/\nu_{2}=\nu_{1}/\nu_{p}$ and $(1+\varepsilon_{p1})(1+\varepsilon_{a2})=1$ holding. In such a case, $\varepsilon_{p1}=(\nu_{1}-\nu_{p\textrm{-sh}})/\nu_{p\textrm{-sh}}$ and $\varepsilon_{a2}=(\nu_{2}-\nu_{a\textrm{-sh}})/\nu_{a\textrm{-sh}}$ have the same $\beta$-solution, because the antiparallel and parallel laser beams interact with the same ion beam.
It should be emphasized that the basic physical requirement for the definition of measurement accuracy of Doppler effect is that Einstein’s Doppler formula must be confirmed when the accuracy is equal to zero. In the single laser beam cases, for example, if $\varepsilon_{p1}=0$, then from Eq. (6) we have $\nu_{1}=\nu_{p\textrm{-sh}}$, and from Eq. (4) we have $\nu_{1}=\gamma(1-\beta)\nu_{p}$, which is the exact Einstein’s Doppler formula for the parallel laser beam for the measured frequency $\nu_{1}$. Putting it simply, Einstein’s Doppler formula $\nu_{1}=\gamma(1-\beta)\nu_{p}$ is strictly confirmed at $\varepsilon_{p1}=0$. Similarly, Einstein’s Doppler formula $\nu_{2}=\gamma(1+\beta)\nu_{a}$ is confirmed at $\varepsilon_{a2}=0$. Thus the definition of the measurement accuracy for a single laser beam is physically justified. However, the definition $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$ in r7 cannot confirm Einstein’s Doppler formula at $\varepsilon=0$, and it is not physical, as shown below.
Inserting Eq. (3) into Eq. (7) yields
$\beta=\frac{\nu^{2}_{2}-\nu^{2}_{a}(1+\varepsilon_{a2})^{2}}{\nu^{2}_{2}+\nu^{2}_{a}(1+\varepsilon_{a2})^{2}}.$ (18)
In the following calculations, the experimental data $\nu_{2}=546,474,960.7$ MHz, $\nu_{a}=384,225,534.98$ MHz ($\lambda_{a}=780$ nm), and $\nu_{p}=777,210,326.98$ MHz ($\lambda_{p}=386$ nm) r7 are taken, while $\nu_{1}=(\nu_{a}\nu_{p})/\nu_{2}$ is assumed to make $\varepsilon=0$ exactly hold.
  1. (i)
    For $\varepsilon=0$, both $\varepsilon_{p1}=0$ and $\varepsilon_{a2}=0$ can hold. In such a case, we have $\nu_{2}=\gamma(1+\beta)\nu_{a}$ (Einstein’s Doppler formula) holding, and we have
    $\beta=\frac{\nu^{2}_{2}-\nu^{2}_{a}}{\nu^{2}_{2}+\nu^{2}_{a}}\approx 0.338,377,220,835.$ (19)
  2. (ii)
    For $\varepsilon=0$, $\varepsilon_{p1}=1$ and $\varepsilon_{a2}=-0.5$ also can hold, not necessarily $\varepsilon_{p1}=0$ and $\varepsilon_{a2}=0$. In such a case we have
    $\beta=\frac{\nu^{2}_{2}-\frac{1}{4}\nu^{2}_{a}}{\nu^{2}_{2}+\frac{1}{4}\nu^{2}_{a}}\approx 0.780,013,866,267,$ (20)
    greatly deviating from $\beta\approx 0.338,377,220,835$ obtained from Einstein’s Doppler formula $\nu_{2}=\gamma(1+\beta)\nu_{a}$ for a single laser beam, as shown in above (i).
Thus from above (i) and (ii), we can conclude that Einstein’s Doppler formula $\nu_{2}=\gamma(1+\beta)\nu_{a}$ is not confirmed at $\varepsilon=0$, because we cannot be sure that $\nu_{2}=\gamma(1+\beta)\nu_{a}$ will hold when $\varepsilon=0$ holds.
In fact, from Eq. (18) one can see that for given $\nu_{2}~{}(\neq 0)$ and $\nu_{a}~{}(\neq 0)$, there are an infinite number of $\beta$-values that make $\varepsilon=0$ hold, where $(-1<\beta<1)$ corresponds to $(+\infty>\varepsilon_{a2}>-1)$. This can also be understood directly from the definition $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$ itself, where $\varepsilon=0$ always holds as long as $\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})=1$ is true, independent of whether $\nu_{i}(i=1,2)$ and $\nu_{j}(j=a,p)$ fulfill Einstein’s Doppler formula.
So far we have finished the proof that, the definition $\varepsilon=\sqrt{\nu_{a}\nu_{p}/(\nu_{1}\nu_{2})}-1$ in r7 indeed does not meet the basic physical requirement for the measurement accuracy of Einstein’s Doppler effect stated previously, and it is not physical.